# Electrical spin manipulation in double SrTiO3/LaAlO3 quantum dots

### v1

@@ III.2 EDSR in weakly asymmetric system
III.3 Optimization of the dipole matrix moment
@@
For the studies of the transition times we chose meV – near the maximum of the dipole matrix element [Fig. 11(a)] and three values of the magnetic field after the singlet-triplet avoided crossing, where the ground-state is the spin triplet – T, T and T – see Fig. 10(b). The singlet-triplet transition times as functions of the AC amplitude is given in Fig. 12. The transition times reach minima for lower values of the AC field amplitude than in Fig. 9(b) and start to grow for the values that exceeds the one for the LZSM transition regime. The LZSM regime is open for the amplitudes of V, 2.7V and 9.5 V/nm for T, T and T, respectively. Note, that for each value of the magnetic field the minimum spin flip time of about 330 ps is achieved for some value of the AC field amplitude. Since the amplitudes of the AC field near the minimal spin flip time are smaller than in Fig. 9(b), so are the wave function leakage values near the minimum which are only 0.5%, 0.02%, 0.0043% for V/nm, V/nm, and V/nm, corresponding to T, T and T, respectively.
III.4 Spin flips for an ideally symmetric system
@@
The study of the AC driven spin-flips in a symmetrix system with nm is summarized in Fig. 14 for two values of the external magnetic field T and T, both corresponding to the ground-state triplet. The position of the resonant second-order transition on the energy scale is redshifted with the amplitude [Fig. 14(a)]. In the studied range of the AC amplitudes – the spin-flip time is a decreasing function of the amplitude [Fig. 14(b)]. For T the LZSM regime opens at 45 eV/nm, and for T at still higher amplitudes which are outside the range studies in Fig. 14.
@@ V ACKNOWLEDGEMENT
This work is financed by the Horizon Europe EIC Pathfinder under the grant IQARO number 101115190 titled "Spin-orbitronic quantum bits in reconfigurable 2D-oxides" and partly supported by program „Excellence initiative – research university” for the AGH University. Computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) within computational grant no. PLG/2023/016317 was used.

### v2

@@ III.2 EDSR in weakly asymmetric system
Due to the very strong spin-orbital interactions in the present system the minimal EDSR-induced slip flip time of 245 ps is very short. Note that the fastest single-triplet transition is obtained when the time evolution is not longer a two-level Rabi cycle with the presence of higher-excited states of about 5% (see Fig. 13(c)). However, even for the lowest considered AC field amplitude, when the spin transitions follow the Rabi cycle, the spin flip times of a few nanoseconds are still very low as compared to the values observed in quantum dots defined in GaAs or Si. The corresponding electron spin inversion time mediated via the spin-orbit coupling in electrostatic double quantum dots in GaAs is 106 ns Nowack et al. (2007). Faster spin-inversion times of about 50 ns have been achieved with the microwave chip generating magnetic field locally in the GaAs quantum dots Koppens et al. (2006) in an electron spin resonance setup. For GaAs quantum dots with the conduction band formed by s-type electron orbitals the spin inversion can also be achieved in the random field of nuclear spins Laird et al. (2009) but the Rabi oscillations in the hyperfine field are very slow with the characteristic time of about 1 microsecond. The effects of the hyperfine interactions should be negligible in the present system due to the d-type of the Ti orbitals that span the space for the confined states. In silicon the hyperfine field can be minimized in isotopically enriched 28Si and the spin-flip time mediated by spin-valley coupling is of the order of 0.5 microsecond Klemt et al. (2023). Advanced quantum registers of spin-qubits in silicon quantum dots are characterized by faster spin-flip times of ns Philips et al. (2022). These devices Philips et al. (2022) are equipped with integrated micromagnets that form a gradient of the magnetic field which mediates the spin flips in external AC electric field that sweeps the confined electrons through the inhomogenous magnetic field region.
III.3 Orientation of the fields and the interdot spacing
In the paper we assume that the magnetic field is oriented perpendicular to the STO/LAO interface and that the AC electric field is oriented parallel to the axis of the double dot system. For a choice of the orientation within the plane of confinement the orbital effects of the magnetic field are absent. The energy spectrum for the weakly asymmetric system of Fig. 5 and the magnetic field oriented parallel to the axis is displayed in Fig. 14(a). One notices that for the in-plane magnetic field orientation the curvature of the energy levels as a function of is missing outside the singlet-triplet avoided-crossing. The avoided crossing is shifted to higher values of the magnetic field by about 0.1T. The spin polarization of the lowest-states as well as the dipole matrix element given in Fig. 14(b) are very similar to the results in the perpendicular magnetic field (Fig. 5(b)) with the exception of a small shift of the maximum due to the shift on the scale of the avoided crossing. The values of the matrix elements, the energy splittings and the spin components are given in Tables IV, V and VI, respectively. Overall, the orientation of the magnetic field turns out to be of a secondary importance.
The AC field oriented in the direction chosen in this paper produces large dipole matrix elements due to the spatial extension of the charge system (see Fig. 1). The field oriented parallel to the axis produces the matrix elements which are by two or three orders of magnitude smaller (Table VII). The present 2D model does not allow for a straightforward incorporation of the AC electric field variation in the direction perpendicular to the plane of confinement. However, as we argued in Ref. Szafran et al. (2024) the EDSR can be understood as a translation of the electron motion induced by AC field to an effective magnetic field that is – as in Rashba type coupling – orthogonal to the direction of the induced electron spatial movement. The effective field needs to be orthogonal to the external magnetic field polarizing the electron spins. Therefore, the AC modulation of the electric potential perpendicular to the plane of confinement should lead to the transitions between the states polarized in this direction. We do not expect the effect to be stronger than for the field oriented along the axis of the dots due to the very strong confinement of the wave functions in the direction (see Fig. 14 of Ref. Szafran et al. (2024)).
The effect of the distance between the Gaussian quantum well centers on the dipole matrix element and the singlet-triplet avoided crossing is given in Fig. 15(a,b). The curves in Fig. 15 are denoted by an integer , where the spacing between the centers is taken as where nm is the lattice constant. The values of nm and 31.2 nm used in the precedent section correspond to and , respectively. The interdot spacing has a pronounced effect on the position of the singlet-triplet avoided crossing that produces the maximum of the dipole matrix element. For a large interdot spacing the singlet and triplet tend to degeneracy that occurs for strictly separated electrons. A small magnetic field promotes the spin-polarized triplet to the ground-state. For a small values of the systems tends to a single elongated quantum dot and a large magnetic field is required for a ground-state spin transition. The maximum of the dipole matrix element is a non-monotonic function of the interdot spacing. In the limit of decoupled quantum dots with absent interdot tunnel coupling the dipole matrix decreases. On the other hand for a small value of the dipole matrix element cannot reach large values due to a limited extension of the wave functions. The magnetic field that is chosen for the spin manipulation needs to be placed off the avoided crossing since the spin states are evenly mixed at the center of the avoided crossing. Note, that the width of the peaks in Fig. 15(a) grows with decreasing which will affect the experimental workpoint.
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | 2.59 | 2.58 | 0.52 | 0.43 |
2 | 0.042 | 0.193 | 0.46 | 0.18 |
3 | ||||
4 | ||||
5 | ||||
6 |
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | 0 | 0 | 0 | 0 |
2 | 0.142 | 0.0598 | 0.0363 | 0.116 |
3 | 0.213 | 0.214 | 0.0245 | 0.329 |
4 | 0.284 | 0.367 | 0.493 | 0.656 |
5 | 1.318 | 1.318 | 1.352 | 1.433 |
6 | 2.640 | 2.640 | 2.674 | 2.755 |
(0.434T) | (0.94T) | (1.5T) | (2T) | |
1 | -1.905 | -1.981 | ||
2 | -1.989 | -1.970 | ||
3 | ||||
4 | 1.992 | 1.192 | 1.992 | 1.991 |
5 | ||||
6 |
0.434T | |
0.94T | |
1.5T | |
2T |
III.4 Optimization of the dipole matrix moment
@@
For the studies of the transition times we chose meV – near the maximum of the dipole matrix element [Fig. 17(a)] and three values of the magnetic field after the singlet-triplet avoided crossing, where the ground-state is the spin triplet – T, T and T – see Fig. 16(b). The singlet-triplet transition times as functions of the AC amplitude is given in Fig. 18. The transition times reach minima for lower values of the AC field amplitude than in Fig. 13(b) and start to grow for the values that exceeds the one for the LZSM transition regime. The LZSM regime is open for the amplitudes of V, 2.7V and 9.5 V/nm for T, T and T, respectively. Note, that for each value of the magnetic field the minimum spin flip time of about 330 ps is achieved for some value of the AC field amplitude. Since the amplitudes of the AC field near the minimal spin flip time are smaller than in Fig. 13(b), so are the wave function leakage values near the minimum which are only 0.5%, 0.02%, 0.0043% for V/nm, V/nm, and V/nm, corresponding to T and T, and T, respectively.
III.5 Spin flips for an ideally symmetric system
@@
The study of the AC driven spin-flips in a symmetric system with nm is summarized in Fig. 20 for two values of the external magnetic field T and T, both corresponding to the ground-state triplet. The position of the resonant second-order transition on the energy scale is redshifted with the amplitude [Fig. 20(a)]. In the studied range of the AC amplitudes – the spin-flip time is a decreasing function of the amplitude [Fig. 20(b)]. For T the LZSM regime opens at 45 eV/nm, and for T at still higher amplitudes which are outside the range studies in Fig. 20.
@@ V ACKNOWLEDGEMENT
This work is financed by the Horizon Europe EIC Pathfinder under the grant IQARO number 101115190 titled "Spin-orbitronic quantum bits in reconfigurable 2D-oxides" and partly supported by program „Excellence initiative – research university” for the AGH University. Computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) within computational grant no. PLG/2024/017175 was used.